# Time Evolution and Thermal Renormalization Group Flow in Cosmology

### v2

@@ Abstract
Time-evolution of the Universe as described by the Friedmann equation can be coupled to equations of motion of matter fields. Quantum effects may be incorporated to improve these classical equations of motion by the renormalization group (RG) running of their couplings. Since temporal and thermal evolutions are linked to each other, astrophysical and cosmological treatements based on zero-temperature RG methods require the extension to finite-temperatures. We propose and explore a modification of the usual finite-temperature RG approach by relating the temperature parameter to the running RG scale as $T\equiv k_{T}=\tau k$ (in natural units), where $k_{T}$ is acting as a running cutoff for thermal fluctuations and the momentum $k$ can be used for the quantum fluctuations. In this approach, the temperature of the expanding Universe is related to the dimensionless quantity $\tau$ (and not to $k_{T}$). We show that by this choice dimensionless RG flow equations have no explicit $k$-dependence, as it is convenient. We also discuss how this modified thermal RG is used to handle high-energy divergences of the RG running of the cosmological constant and to ”solve the triviality”of the $\phi^{4}$ model by a thermal phase transition in terms of $\tau$ in $d=4$ Euclidean dimensions.
@@ I Introduction
There is an increasing interest in astrophysics and cosmology to incorporate quantum effects by renormalization group (RG) running wilson such as in the asymptotically safe gravity approach reuter , for a recent review see e.g. reuter_saueressig . Application to cosmology bonanno_saueressig ; alessia ; rg_cosmo_12 ; rg_cosmo_17 ; rg_cosmo_18 and to black-hole physics have been extensively discussed cosmo_qeg ; qeg8 . In these approaches, the running RG cutoff $k$ is identified with a typical energy scale of the system cosmo_qeg and the RG running is used to connect the physics of different energy scales and at different times. The RG running of the parameters (couplings) is then incorporated into the classical equations of motion, i.e., the Friedmann equations describing the time-evolution of a homogeneous and isotropic universe.
@@
Finite temperature formalism requires an introduction of a ”finite volume” for the (imaginary) time integral. However, if one follows the Wilsonian approach, this ”finite volume” has to be rescaled in every blocking step, which means it has to be linked to the running momentum cutoff $k$. Since in the RG approach the infrared (IR) limit $k\to 0$ has to be taken, this ”finite volume” cannot be considered as the (inverse) temperature, but it can rather be used as a running momentum cutoff: $T\equiv k_{T}$. Thus, we suggest and explore in the following the identification
@@ II Thermal RG equation
In general, RG allows to treat the effects of fluctuations via the successive elimination of the degrees of freedom which lie above a running momentum cutoff $k$ and generates the momentum-shell functional RG flow equation eea_rg . In the finite-temperature formalism, momentum integrals of the zero-temperature RG equation eea_rg are modified by using Matsubara frequency summation
@@
which is a summation over discrete imaginary frequencies – for bosonic degrees of freedom: $\omega_{n}=2\pi nT$. Typically, one assumes that the regulator function $R_{k}(p)$ of the zero-temperature RG method is independent of the Matsubara frequency, but otherwise the same as for the zero-temperature case, where a usual choice is the Litim reglator $R_{k}(p)=(k^{2}-p^{2})\Theta(k^{2}-p^{2})$ Litim2000 . In the local potential approximation (LPA) when one does not consider any wave function renormalization and the couplings of the scaling potential $V_{k}(\phi)$ carry the RG scaling, the RG equation at finite temperature $T$ with the (frequency-independent) Litim regulator is written as thermal_rg_phi4 :
@@
The thermal RG (4) has a singularity structure identical to its zero-temperature counterpart and determined by $k^{2}+\partial^{2}_{\phi}V_{k}=0$. In Appendix A we show that the dimensionless RG flow equations derived from (4) have no explicit $k$-dependence. This is not the case if $T$ is fixed, i.e., not $k$-dependent. This implies that the choice (1) leads to non-trivial fixed points, unlike the choice of a non-running temperature. This observation is one of the main results of the paper.
@@ III Inflationary Cosmology
The observation of the Higgs boson renewed research activity where the inflaton is associated with the SM Higgs field higgs_inf_1 ; higgs_inf_2 ; higgs_inf_3 ; higgs_inf_4 ; higgs_inf_5 introduces a large non-minimal coupling $\xi$ between the Higgs boson and the Ricci curvature scalar. In case of non-minimal coupling, the Einstein frame where slow-roll study is performed and the Jordan-frame where the RG flow is considered are related to each other via a non-linear transformation. RG transformations generate an additional, $R^{2}$ term. However, we do not consider here quantum effects for gravity, only for the scalar matter field, and thus we choose
@@
$V(\phi=0)\equiv m_{p}^{2}\Lambda,$ (5)
where $\Lambda$ denotes the cosmological constant and not the UV momentum cutoff. We assume an expanding homogeneous and isotropic Universe (with a flat curvature), and with the Friedmann–Lemaître–Robertson–Walker metric, $\sqrt{-g}=\sqrt{-\det(g_{\mu\nu})}=a^{3}$ where the time-dependence of the scalar factor $a(t)$ is given by the Friedmann equation coupled to the equation of motion of the scalar field. This scenario is the an ”economical” one for the theoretical framework for slow-roll inflation where the usual form for the Higgs-like potential is
@@
which is the SM symmetry breaking potential where $v$ is the vacuum expectation value (VeV) and $g_{4}$ is the quartic coupling (dimensionless in $d=4$ dimensions). The slow-roll study of (6) has to confronted to observations, e.g., to most recent Planck data planck on CMBR anisotropy. If one assumes a non-minimal coupling to gravity, i.e., $\xi\neq 0$, the potential has to be written in the Einstein frame to perform the slow-roll study. This transformation is non-linear and have a simplified form for large non-minimal coupling ($\xi\gg 1$): $\varphi\approx m_{p}\sqrt{\frac{3}{2}}\ln(F)$ and $U(\varphi)\equiv m_{p}^{2}V(\phi)/F^{2}(\phi)$ where $F(\phi)=1+\xi\phi^{2}/m_{p}^{2}$. The Higgs-like potential (6) redas in the Einstein frame:
@@ IV Thermal RG and Triviality in Cosmology
where dimensionless quantities denoted by the tilde superscript. Thermal RG flow equations for the first three (dimensionful) couplings are given in Appendix A where we show that the dimensionless RG flow equations have no explicit k-dependence and discuss also the subtraction method rg_cosmo_const for the field-independent couplig $g_{0,k}$ which is related to the cosmological constant.
It is well-known that the Wilson-Fisher fixed point of the $\phi^{4}$ model coincides with the Gaussian one in $d=4$ dimensions. Although the model has two phases, RG trajectories bifurcate from the vertical line at the vanishing mass (when $\tilde{g}_{2,k}=0$), see solid lines on Fig. 1.
Figure 1: Thermal RG flow of the $\phi^{4}$ model in $d=4$ dimensions. Dashed lines show how the separatrix changes with $\tau$.
The classical analysis gives exactly the same result: If the squared mass is negative the potential has a double-well structure which signals the broken phase where the $Z_{2}$ reflection symmetry is broken spontaneously. If the squared mass is positive the potential has a single minimum which is the symmetric phase. The numerical solution of the (dimensionless) thermal RG flow equations (see Appendix A) modifies the flow diagram. With non-vanishing value for $\tau$, the RG trajectory which separates the phases is no longer a vertical line, see dashed lines on Fig. 1. For increasing values of $\tau$, the broken phase is decreased and for $\tau\to\infty$ it is not possible to find any starting point in the vicinity of the Gaussian fixed point from which an RG trajectory can run into the broken phase. Since $\tau$ measures how thermal fluctuations are important compared to quantum fluctuations, one can say that by changing $\tau$ a thermal phase transition occurs. One can always determine a critical value $\tau_{c}$ for any starting point close to the Gaussian fixed point, see the black cross on Fig. 2. If $\tau<\tau_{c}$ the RG trajectory from that starting point runs into the broken (low-temperature) phase and for $\tau>\tau_{c}$ the RG trajectory ends up in the symmetric (high-temperature) phase. Thus, the RG flow diagram is no longer ”trivial” and a thermal phase transition is observed in terms of $\tau$.
@@ V Conclusions
A modification of the usual finite-temperature renormalization (RG) approach is proposed by relating the temperature parameter to the running RG scale: $T\equiv k_{T}=\tau k$. This choice is convenient since with it the dimensionless RG equation has no explicit $k$-dependence, see Appendix A, and this is crucial to have fixed points. We applied this thermal RG to handle UV divergences in the RG running of the cosmological constant, see Appendix A, and to solve the triviality $\phi^{4}$ theory in $d=4$. These results suggest to associate the parameter $\tau$ with the temperature of the expanding Universe, so $g_{2}=g_{2}(\tau)$ and $g_{4}=g_{4}(\tau)$ are assumed. The quantum effects are taken into account by the thermal RG running of the couplings at a given temperature, i.e., one has to take the limit $k\to 0$ by moving along the thermal RG trajectory with fixed value of $\tau$ with given initial conditions. Then $\tau$ can be decreased and the thermal RG procedure repeated, producing the quantum effective action at a given temperature. Time and thermal evolutions of the Universe are linked to each other, so the change in $\tau$ connects the scales of the cosmic inflation and the electroweak phase transition and thermal phase transition of the $\phi^{4}$ theory validates the vanishing VeV in Eq. (6) when the slow-roll study of Higgs inflation is performed by Eq. (7). Finally, we observe that since perturbative and non-perturbative RG equations are related to each other scheme1 ; scheme2 ; scheme3 ; frg_prg , one expects that our modified thermal RG method with $T\equiv k_{T}=\tau k$ finds application in particle cosmology both in perturbative and non-perturbative approaches.
Acknowledgements.
@@
Appendix A Thermal RG flow equations the of the $\phi^{4}$ model in $d=4$ Euclidean dimensions.
Although particle physics and cosmology are intimately connected, in accelerators a limited number of colliding particles are taken into account while the early Universe must be seen as a hot dense plasma. Thus, the zero temperature quantum field theory which is the usual theoretical framework for particle processes must be extended to finite temperatures for cosmological applications. Indeed, thermal field theory which is a very well-developed framework for many-body physics negele ; kapusta relates a quantum system in $d-1$ (spatial) dimensions at $T=0$ and the corresponding classic system in $d$ dimensions at finite temperature sachdev .
However, there are open question on its connection to renormalization and renormalization group (RG) applications. For example, in the perturbative RG approach the temperature $T$ is fixed to the running RG momentum scale $\mu=2\pi T$ (in natural units), while in non-perturbative RG it is linked to the the ultraviolet (UV) initial value $\Lambda$ of the running momentum cutoff by putting $T=\tau\Lambda$ (in natural units), see e.g., thermal_rg_phi4 ; thermal_rg_phi4_pressure ; thermal_rg_phi4_volume . The latter choice is motivated by the fact that the RG scale-dependence is introduced artificially in the Wilsonian approach wilson and the quantized theory must be obtained in the physical limit where the RG scale $k$ is sent to zero, i.e., $k\to 0$, and it seems reasonable to connect the temperature to a fixed (i.e, not running) momentum. However, it has a serious drawback: one finds an explicit RG scale dependence in the dimensionless RG flow equations which makes no room for non-trivial fixed points.
We propose, here a modification of the usual finite-temperature RG approach by relating the temperature parameter to the running RG scale by a variable dimensionless factor $\tau$ which plays the role of the temperature of the cosmic plasma, i.e, $T\equiv k_{T}=\tau k$ (in natural units). In this case, the parameter $T\equiv k_{T}$ annot be considered as the real temperature but rather be used as a running momentum cutoff while the temperature of the expanding Universe is related to the dimensionless quantity $\tau$. By this choice we show that the dimensionless RG flow equations have no explicit RG scale-dependence.
Indeed, by using $T\equiv k_{T}=\tau k$ the following thermal RG equation can be written, see also (4)
@@
$\left(d-\frac{(d-2)}{2}\tilde{\phi}\partial_{\tilde{\phi}}+k\partial_{k}\right)$ $\tilde{V}_{k}(\tilde{\phi})=\frac{2\alpha_{d-1}}{d-1}\,\frac{\coth{\left(\frac{\sqrt{1+\partial^{2}_{\tilde{\phi}}\tilde{V}_{k}(\tilde{\phi})}}{2\tau}\right)}}{2\sqrt{1+\partial^{2}_{\tilde{\phi}}\tilde{V}_{k}(\tilde{\phi})}}\,,$ (9)
@@
$\left(d-\frac{(d-2)}{2}\tilde{\phi}\partial_{\tilde{\phi}}+k\partial_{k}\right)$ $\tilde{V}_{k}(\tilde{\phi})=\frac{2\alpha_{d-1}}{d-1}\,\frac{\coth{\left(\frac{k}{\Lambda}\frac{\sqrt{1+\partial^{2}_{\tilde{\phi}}\tilde{V}_{k}(\tilde{\phi})}}{2\tau}\right)}}{2\sqrt{1+\partial^{2}_{\tilde{\phi}}\tilde{V}_{k}(\tilde{\phi})}}\,,$ (10)
which makes no room for non-trivial fixed point solutions, i.e., by setting $k\partial_{k}\tilde{V}_{k}\equiv 0$, the remaining algebric equation can have only k-dependent solution.
@@
$k\partial_{k}$ $g_{2,k}=\frac{1}{6\pi^{2}}k^{5}\left[-\frac{g_{4,k}\coth\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}{4\left(k^{2}+g_{2,k}\right)^{3/2}}-\frac{g_{4,k}\;\text{csch}^{2}\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}{8\tau k\left(k^{2}+g_{2,k}\right)}\right],$ (12)
$k\partial_{k}$ $g_{4,k}=\frac{1}{6\pi^{2}}k^{5}\left[\frac{9g_{4,k}^{2}\coth\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}{8(k^{2}+g_{2,k})^{5/2}}+\frac{9g_{4,k}^{2}\text{csch}^{2}\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}{16\tau k(k^{2}+g_{2,k})^{2}}+\frac{3g_{4,k}^{2}\coth\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)\text{csch}^{2}\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}{16\tau^{2}k^{2}(k^{2}+g_{2,k})^{3/2}}\right]$ (13)
@@
$k\partial_{k}g_{0,k}$ $=$ $\frac{1}{6\pi^{2}}k^{5}\frac{\coth{\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}}{2\sqrt{k^{2}+g_{2,k}}}\frac{k^{4}}{12\pi^{2}}\coth{\frac{1}{2\tau}}-\frac{k^{2}}{24\pi^{2}}\frac{1+\tau\sinh{\frac{1}{\tau}}}{\tau-\tau\cosh{\frac{1}{\tau}}}g_{2,k}\,,$ (14)
where the subtracted terms do not violate the IR behaviour since they vanish for $k\to 0$. The RG flow for the cosmological constant can be derived from (14) based on the relation
$V(\phi=0)\equiv m_{p}^{2}\Lambda,$ (15)
where $\Lambda$ denotes the cosmological constant and not the UV momentum cutoff, which gives

### v3

@@ Abstract
Time-evolution of the Universe as described by the Friedmann equation can be coupled to equations of motion of matter fields. Quantum effects may be incorporated to improve these classical equations of motion by the renormalization group (RG) running of their couplings. Since temporal and thermal evolutions are linked to each other, astrophysical and cosmological treatments based on zero-temperature RG methods require the extension to finite-temperatures. We propose and explore a modification of the usual finite-temperature RG approach by relating the temperature parameter to the running RG scale as $T\equiv k_{T}=\tau k$ (in natural units), where $k_{T}$ is acting as a running cutoff for thermal fluctuations and the momentum $k$ can be used for the quantum fluctuations. In this approach, the temperature of the expanding Universe is related to the dimensionless quantity $\tau$ (and not to $k_{T}$). We show that by this choice dimensionless RG flow equations have no explicit $k$-dependence, as it is convenient. We also discuss how this modified thermal RG is used to handle high-energy divergences of the RG running of the cosmological constant and to ”solve the triviality”of the $\phi^{4}$ model by a thermal phase transition in terms of $\tau$ in $d=4$ Euclidean dimensions.
@@ I Introduction
There is an increasing interest in astrophysics and cosmology to incorporate quantum effects by renormalization group (RG) running wilson such as in the asymptotically safe gravity approach reuter , for a recent review see e.g. reuter_saueressig ; AS_grav_1 ; AS_grav_2 . Application to cosmology bonanno_saueressig ; alessia ; rg_cosmo_12 ; rg_cosmo_17 ; rg_cosmo_18 and to black-hole physics have been extensively discussed cosmo_qeg ; qeg8 . In these approaches, the running RG cutoff $k$ is identified with a typical energy scale of the system cosmo_qeg and the RG running is used to connect the physics of different energy scales and at different times. The RG running of the parameters (couplings) is then incorporated into the classical equations of motion, i.e., the Friedmann equations describing the time-evolution of a homogeneous and isotropic universe.
@@
Finite temperature formalism requires an introduction of a ”finite volume” for the (imaginary) time integral. However, if one follows the Wilsonian approach, this ”finite volume” has to be rescaled in every blocking step, which means it has to be linked to the running momentum cutoff $k$. Since in the RG approach the infrared (IR) limit $k\to 0$ has to be taken, this ”finite volume” cannot be considered as the (inverse) temperature, but it can rather be used as a running momentum cutoff: $T\equiv k_{T}$. Thus, we suggest and explore the following identification
@@
It is important to note that all computations of the present paper are performed in the Euclidean spacetime. In order to connect the results with cosmology and particle physics, one could speculate that there is the need for justification of the use of Euclidean metric. Indeed, there is an increasing interest in the literature to discuss whether an analytic continuation to Minkowski spacetime is possible in the framework of the Wilsonian RG method, see e.g., minkowski_frg . However, let us note that lattice calculations are also performed in the Euclidean spacetime with huge amount of phenomenological applications, which serve as a strong support for the use of Euclidean instead of the Minkowski spacetime.
@@ II Thermal RG equation
The RG concept plays an important role in the description of physical systems across different scales. The Wilsonian RG method has been constructed to perform the renormalization non-perturbatively wilson . In general, RG allows to treat the effects of fluctuations via the successive elimination of the degrees of freedom which lie above a running momentum cutoff $k$ and generates the momentum-shell functional RG flow equation, i.e., the Wetterich equation eea_rg . At zero-temperature it is formulated in Euclidean spacetime and stands for the running effective action $\Gamma_{k}[\phi]$ with its Hessian $\Gamma^{(2)}_{k}[\phi]$ where $k$ is the running momentum cutoff (i.e., the RG scale). The Wetterich equation eea_rg contains, an appropriately chosen regulator function $R_{k}$, which is introduced to decouple slow modes with low momenta while leaving high momentum modes unaffected. The regulator ensures that the effective action captures all relevant fluctuations.
In the finite-temperature formalism, momentum integrals of the zero-temperature RG equation eea_rg are modified by using Matsubara frequency summation
@@
which is a summation over discrete imaginary frequencies – for bosonic degrees of freedom: $\omega_{n}=2\pi nT$. Typically, one assumes that the regulator function $R_{k}(p)$ of the zero-temperature RG method is independent of the Matsubara frequency, but otherwise the same as for the zero-temperature case, where a usual choice is the Litim regulator $R_{k}(p)=(k^{2}-p^{2})\Theta(k^{2}-p^{2})$ Litim2000 . In the local potential approximation (LPA) when one does not consider any wave function renormalization and the couplings of the scaling potential $V_{k}(\phi)$ carry the RG scaling, the RG equation at finite temperature $T$ with the (frequency-independent) Litim regulator is written as thermal_rg_phi4 :
@@
The thermal RG (4) has a singularity structure identical to its zero-temperature counterpart and determined by $k^{2}+\partial^{2}_{\phi}V_{k}=0$. In  A we show that the dimensionless RG flow equations derived from (4) have no explicit $k$-dependence. This is not the case if $T$ is fixed, i.e., not $k$-dependent. This implies that the choice (1) can lead to non-trivial fixed points, unlike the choice of a non-running temperature. This observation is one of the main results of the paper.
@@ III Inflationary Cosmology
The observation of the Higgs boson renewed research activity where the inflaton is associated with the SM Higgs field higgs_inf_1 ; higgs_inf_2 ; higgs_inf_3 ; higgs_inf_4 ; higgs_inf_5 introduces a large non-minimal coupling $\xi$ between the Higgs boson and the Ricci curvature scalar higgs_inf_6 ; higgs_inf_7 ; higgs_inf_8 ; higgs_inf_9 . In case of non-minimal coupling, the Einstein frame where slow-roll study is performed and the Jordan-frame where the RG flow is considered are related to each other via a non-linear transformation. RG transformations generate an additional, $R^{2}$ term. However, we do not consider here quantum effects for gravity, only for the scalar matter field, and thus we choose
@@
$V(\phi=0)\equiv m_{p}^{2}\Lambda_{{\mathrm{cosmo}}},$ (5)
where $\Lambda_{{\mathrm{cosmo}}}$ denotes the cosmological constant and not the UV momentum cutoff. We assume an expanding homogeneous and isotropic Universe (with a flat curvature), and with the Friedmann–Lemaître–Robertson–Walker metric, $\sqrt{-g}=\sqrt{-\det(g_{\mu\nu})}=a^{3}$ where the time-dependence of the scalar factor $a(t)$ is given by the Friedmann equation coupled to the equation of motion of the scalar field. This scenario is the an ”economical” one for the theoretical framework for slow-roll inflation where the usual form for the Higgs-like potential is
@@
which is the SM symmetry breaking potential where $v$ is the vacuum expectation value (VeV) and $g_{4}$ is the quartic coupling (dimensionless in $d=4$ dimensions). The slow-roll study of (6) has to confronted to observations, e.g., to most recent Planck data planck on CMBR anisotropy. If one assumes a non-minimal coupling to gravity, i.e., $\xi\neq 0$, the potential has to be written in the Einstein frame to perform the slow-roll study. This transformation is non-linear and have a simplified form for large non-minimal coupling ($\xi\gg 1$): $\varphi\approx m_{p}\sqrt{\frac{3}{2}}\ln(F)$ and $U(\varphi)\equiv m_{p}^{2}V(\phi)/F^{2}(\phi)$ where $F(\phi)=1+\xi\phi^{2}/m_{p}^{2}$. The Higgs-like potential (6) reads in the Einstein frame:
@@ IV Thermal RG and Triviality in Cosmology
where dimensionless quantities denoted by the tilde superscript. Thermal RG flow equations for the first three (dimensionful) couplings are given in  A where we show that the dimensionless RG flow equations have no explicit k-dependence and discuss also the subtraction method rg_cosmo_const for the field-independent coupling $g_{0,k}$ which is related to the cosmological constant.
It is well-known that the Wilson-Fisher fixed point of the $\phi^{4}$ model coincides with the Gaussian one in $d=4$ dimensions vbf_mati . Although the model has two phases, RG trajectories bifurcate from the vertical line at the vanishing mass (when $\tilde{g}_{2,k}=0$), see solid lines on Fig. 1.
Figure 1: Thermal RG flow of the $\phi^{4}$ model in $d=4$ dimensions. Dashed lines show how the separatrix changes with $\tau$. The separatrix belongs to $\tau=0$ (most right dashed line) almost overlaps with the vertical axis (thin solid line).
The classical analysis gives exactly the same result: If the squared mass is negative the potential has a double-well structure which signals the broken phase where the $Z_{2}$ reflection symmetry is broken spontaneously. If the squared mass is positive the potential has a single minimum which is the symmetric phase. The numerical solution of the (dimensionless) thermal RG flow equations (see  A) modifies the flow diagram. With non-vanishing value for $\tau$, the RG trajectory which separates the phases is no longer a vertical line, see dashed lines on Fig. 1. For increasing values of $\tau$, the broken phase is decreased and for $\tau\to\infty$ it is not possible to find any starting point in the vicinity of the Gaussian fixed point from which an RG trajectory can run into the broken phase. Since $\tau$ measures how thermal fluctuations are important compared to quantum fluctuations, one can say that by changing $\tau$ a thermal phase transition occurs. One can always determine a critical value $\tau_{c}$ for any starting point close to the Gaussian fixed point, see the black cross on Fig. 2. If $\tau<\tau_{c}$ the RG trajectory from that starting point runs into the broken (low-temperature) phase and for $\tau>\tau_{c}$ the RG trajectory ends up in the symmetric (high-temperature) phase. Thus, the RG flow diagram is no longer ”trivial” and a thermal phase transition is observed in terms of $\tau$.
@@ V Conclusions
A modification of the usual finite-temperature renormalization (RG) approach is proposed by relating the temperature parameter to the running RG scale: $T\equiv k_{T}=\tau k$. This choice is convenient since with it the dimensionless RG equation has no explicit $k$-dependence, see  A, and this is crucial to have fixed points. We applied this thermal RG to handle UV divergences in the RG running of the cosmological constant, see  A, and to solve the triviality $\phi^{4}$ theory in $d=4$. These results suggest to associate the parameter $\tau$ with the temperature of the expanding Universe, so $g_{2}=g_{2}(\tau)$ and $g_{4}=g_{4}(\tau)$ are assumed. The quantum effects are taken into account by the thermal RG running of the couplings at a given temperature, i.e., one has to take the limit $k\to 0$ by moving along the thermal RG trajectory with fixed value of $\tau$ with given initial conditions. Then $\tau$ can be decreased and the thermal RG procedure repeated, producing the quantum effective action at a given temperature. Time and thermal evolutions of the Universe are linked to each other, so the change in $\tau$ connects the scales of the cosmic inflation and the electroweak phase transition and thermal phase transition of the $\phi^{4}$ theory validates the vanishing VeV in Eq. (6) when the slow-roll study of Higgs inflation is performed by Eq. (7).
We observe that since perturbative and non-perturbative RG equations are related to each other scheme1 ; scheme2 ; scheme3 ; frg_prg , one expects that our modified thermal RG method with $T\equiv k_{T}=\tau k$ finds application in particle cosmology both in perturbative and non-perturbative approaches. However, let us note that it is a valid question to ask whether the Wilsonian RG running scale parameter $k$ can or cannot be identified with physical properties of the system. In order to compute observables using the effective action one has to take the physical limit $k\to 0$ which is not yet possible in reasonable approximations. Although, there are RG studies of cosmological problems where the Wilsonian flow with the scale ”k” is used as an approximation to the physical flow with respect to the physical momentum ”p”, in this letter we relay on the effective action taken in the physical limit, i.e., $k\to 0$ form where physical quantities can be extracted. In other words, we do not consider the k-flow as an approximation to catch the main physics properties of the system at a given range of energy.
Finally, let us mention that an interesting direction of research would be to discuss $T\to\infty$ considering quantum gravity effects that should become relevant at the Planck scale, in order to understand how does the inclusion of gravity affects the proposed scheme. It would be also useful to apply the new thermal RG approach proposed here with the introduction of the dimensionless temperature $\tau$ to investigate in this context the QPT-CPT phase diagram of quantum field theories qpt_cpt_thermal_rg , such as the $\phi^{4}$ and the sine-Gordon models in lower dimensions. The so called QPT-CPT diagram represents graphically the interplay between the classical phase transitions (CPT) and the quantum phase transitions (QPT) which has also been investigated in connection to the Naturalness/Hierarchy problem BrBrCoDa ; BrBrCo ; qpt_higgs1 ; qpt_higgs2 . For example, in qpt_higgs1 ; qpt_higgs2 its was shown that the hierarchy problem as well as the metastability of the electroweak vacuum can be understood as the Higgs potential being near-critical, i.e., close to a QPT.
Acknowledgements
@@
Appendix A Thermal RG flow equations of the $\phi^{4}$ model in $d=4$ Euclidean dimensions.
By using $T\equiv k_{T}=\tau k$ the following thermal RG equation can be written, see also (4)
@@
$\left(d-\frac{(d-2)}{2}\tilde{\phi}\partial_{\tilde{\phi}}+k\partial_{k}\right)$ $\tilde{V}_{k}(\tilde{\phi})=$ (9)
$\frac{2\alpha_{d-1}}{d-1}\,$ $\frac{\coth{\left(\frac{\sqrt{1+\partial^{2}_{\tilde{\phi}}\tilde{V}_{k}(\tilde{\phi})}}{2\tau}\right)}}{2\sqrt{1+\partial^{2}_{\tilde{\phi}}\tilde{V}_{k}(\tilde{\phi})}}\,,$
@@
$\left(d-\frac{(d-2)}{2}\tilde{\phi}\partial_{\tilde{\phi}}+k\partial_{k}\right)$ $\tilde{V}_{k}(\tilde{\phi})=$ (10)
$\frac{2\alpha_{d-1}}{d-1}\,$ $\frac{\coth{\left(\frac{k}{\Lambda}\frac{\sqrt{1+\partial^{2}_{\tilde{\phi}}\tilde{V}_{k}(\tilde{\phi})}}{2\tau}\right)}}{2\sqrt{1+\partial^{2}_{\tilde{\phi}}\tilde{V}_{k}(\tilde{\phi})}}\,,$
which makes no room for non-trivial fixed point solutions, i.e., by setting $k\partial_{k}\tilde{V}_{k}\equiv 0$, the remaining algebraic equation can have only k-dependent solution.
@@
$k\partial_{k}g_{2,k}=$ $\frac{1}{6\pi^{2}}k^{5}\left[-\frac{g_{4,k}\coth\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}{4\left(k^{2}+g_{2,k}\right)^{3/2}}\right.$ (12)
$\left.-\frac{g_{4,k}\;\text{csch}^{2}\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}{8\tau k\left(k^{2}+g_{2,k}\right)}\right],$
$k\partial_{k}g_{4,k}=$ $\frac{1}{6\pi^{2}}k^{5}\left[\frac{9g_{4,k}^{2}\coth\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}{8(k^{2}+g_{2,k})^{5/2}}\right.$ (13)
$\left.+\frac{9g_{4,k}^{2}\text{csch}^{2}\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}{16\tau k(k^{2}+g_{2,k})^{2}}\right.$
$\left.+\frac{3g_{4,k}^{2}\coth\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)\text{csch}^{2}\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}{16\tau^{2}k^{2}(k^{2}+g_{2,k})^{3/2}}\right]$
@@
$k\partial_{k}g_{0,k}=$ $\frac{1}{6\pi^{2}}k^{5}\frac{\coth{\left(\frac{\sqrt{k^{2}+g_{2,k}}}{2\tau k}\right)}}{2\sqrt{k^{2}+g_{2,k}}}-\frac{k^{4}}{12\pi^{2}}\coth{\frac{1}{2\tau}}$
$-\frac{k^{2}}{24\pi^{2}}\frac{1+\tau\sinh{\frac{1}{\tau}}}{\tau-\tau\cosh{\frac{1}{\tau}}}g_{2,k}\,,$ (14)
where the subtracted terms do not violate the IR behaviour since they vanish for $k\to 0$. The RG flow for the cosmological constant can be derived from (A) based on the relation
$V(\phi=0)\equiv m_{p}^{2}\Lambda_{{\mathrm{cosmo}}},$ (15)
where $\Lambda_{{\mathrm{cosmo}}}$ denotes the cosmological constant and not the UV momentum cutoff, which gives